OpenStudy (toxicsugar22):

4x^2-6y^2+32x-24y+16=0 give the exact coordinated of an four points

OpenStudy (amistre64):

(4x^2 +32x +____) -(6y^2 -24y + ____) = 16 +4(__)-6(___) 4(x^2 +8x +16) -6(y^2 -4y +4) = 16 +4(16)-6(4)

OpenStudy (amistre64):

4(x+4)^2 -6(y-2)^2 = 16 +64-24 4(x+4)^2 -6(y-2)^2 = 24 4/24(x+4)^2 -6/24(y-2)^2 = 1 (1/6)(x+4)^2 -(1/4)(y-2)^2 = 1

OpenStudy (amistre64):

when y=0; (x+4)^2 = 12 x+4 = +- 4sqrt(3) x = -4 +- 4sqrt(3) (-4 + 4sqrt(3),0) (-4 + 4sqrt(3),0) are 2 points

OpenStudy (amistre64):

when y = 2; (x+4)^2 = 6 x = -4 +- sqrt(6) (-4 + sqrt(6),2) (-4 -sqrt(6),2) are 2 mor epoints

OpenStudy (toxicsugar22):

thanks